X^3+2 x^2+3 x+6 = 0 rational root
Answer by stanbon(75887)

X^3 + 2x^2 + 3x + 6 = 0 rational root
---------
I graphed it and found a root at x = -2
---
Use synthetic division to find the other:
-2)....1....2....3....6
........1.....0...3...|..0

Quotient: x^2 + 3
Remainder: 0
=======================
Solve x^2+3 = 0
x = sqrt(3) or x = -sqrt(3)
===============================
Answer: x = -2, sqrt(3), -sqrt(3)
===============================
Cheers,
Stan H.